OpenStudy (anonymous):

find the 8th term of each arithmetic sequence. a1=3, d=2

OpenStudy (anonymous):

\[a _{n}=a1(n-1)d\] \[a _{n}=3(8-1)2\] \[3(7)2\] \[3(14)\]

OpenStudy (anonymous):

think you can solve from there